OpenStudy (anonymous):

Solve the equation 2 over 5 (x − 4) = 2x for x. −8 negative 1 over 2 1 −1

OpenStudy (anonymous):

will give medal

OpenStudy (theviper):

Can u draw it ? & to call anyone for help just mention him such as @Para_phrenic88 :)

OpenStudy (theviper):

Is it like this ? \[\Large{\frac{2}{5x-4}=2x}\]

OpenStudy (anonymous):

-8?

OpenStudy (theviper):

If yes then ans will be\[\Large{\frac{2}{5x-4}=2x}\]\[\large{\implies2=2x(5x-4)}\]\[\large{\implies2=10x^2-8x}\]\[\large{\implies10x^2-8x-2=0}\]Now we take 2 as common\[\large{2(5x^2-4x-1)=0}\]\[\large{\implies5x^2-4x-1=0}\]Now solve it by Quadratic Formula :)

OpenStudy (anonymous):

ugh i suck at algebra :(